5a-10b=20
10a+5b=15

Answer by mananth(16949)

5 a -10 b = 20 .............1
10 a + 5 b = 15 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 2
5 a -10 b = 20
20 a + 10 b = 30
Add the two equations
25 a = 50
/ 25
a = 2
plug value of a in (1)
5 a + -10 b = 20
10 + -10 b = 20
-10 b = 20 -10
-10 b = 10
b = -1